# Sub-shot noise sensitivity via deformed four-headed kitten states

### v1

@@ III.5 Enhancement in Sensitivity
Sensitivity and its relationship to the phase-space characteristics of a quantum state are thoroughly discussed in the §II.1. This concept is then applied to the compass state in the §II.2, where it is demonstrated that the presence of sub-Planck structures in those states significantly have enhanced their sensitivity to displacement and that optimizing controlling parameters may further enhance this sensitivity far better than the standard quantum limit. In §III.4, we thoroughly examined the phase space of the proposed SA and AS cases, confirming the presence of sub-Planck structures in their phase spaces. We will now examine how these sub-Planck structures have impact on the sensitivity to phase-space displacement, which is analyzed by assessing the sensitivities using Eq. (3).
@@
$O_{\ket{\square_{i}}\bra{\square_{j}}}(\delta):=$ $G_{\alpha_{i},\alpha_{j}}\Lambda\sum^{p}_{n=0}(-1)^{n}\Gamma H_{p-n,q}\left[\text{i}\big{(}\alpha^{*}_{i}-\delta^{*}\big{)},\text{i}\alpha_{j}\right]$
×$\times H_{p-n,q}\left[-\text{i}\big{(}\alpha_{j}+\delta\big{)},-\text{i}\alpha^{*}_{i}\right]$ (34)
@@
$O_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\delta):=$ $G_{\alpha_{i},\alpha_{j}}\Lambda\sum^{p}_{n=0}(-1)^{n}\Gamma H_{p-n,q}\left[\alpha^{*}_{i}-\delta^{*},\alpha_{j}\right]$
×$\times H_{p-n,q}\left[\alpha_{j}+\delta,\alpha^{*}_{i}\right].$ (37)
@@
The corresponding sensitivities $S_{\ket{\square}}(\delta)$ and $S_{\ket{\blacksquare}}(\delta)$ with $\delta:=\nicefrac{{(\delta x+\text{i}\delta p)}}{{\sqrt{2}}}$ are shown in the Figs. 8(a)-(c) and Figs. 8(d)-(f), respectively.
@@
Let us now examine the SA case presented in the Figs. 8(a)-(c), where it is evident that the overlap is zero for values of $|\delta|<1$ along any direction in phase space. This indicates that the sensitivity to displacement in this scenario surpasses the standard limit. The enhancement in sensitivity becomes more pronounced as the parameter $p$, which represents the number of added photons to the kitten state, increases. As shown in Figs. 8(a) and 8(b), this effect becomes clearly noticeable with a higher number of added photons. Specifically, the central structure is significantly reduced when the number of added photons increases from $p=12$ to $p=24$ with the number of subtracted photons, $q=12$, remains constant in this situation. Furthermore, the impact of increasing the number of eliminated photons from $q=12$ to $q=20$ is illustrated in the Figs. 8(b) and 8(c), where an enlargement in the central structure occurs, indicating that a higher value of $|\delta|$ relative to the earlier case represented in Fig. 8(b) is needed to make the overlap zero. This suggests that the sensitivity to displacement in this circumstance decreases as $q$ grows; hence, contrary to $p$, increasing the number of subtracted photons reduces sensitivity in the SA scenario.
@@
In the AS instance, as illustrated in Figs. 8(d)-(f), similar to the SA case, an improved sensitivity to phase space displacements $\delta$ is noticed, and the values $|\delta|<1$ can render the $S_{\ket{\blacksquare}}(\delta)$ zero as observable by the central structure. Similar to the SA example, this amplification becomes more noticeable as $p$ increases, as illustrated in the instances depicted in Figs. 8(d) and 8(e), and then when $q$ is increased, the improvement in the sensitivity is reduced. Isotropic regions, such as the circles centered at the origin shown in Figs. 8(d) and 8(e), indicates that in the AS case, sensitivity is uniformly enhanced in all directions.
In summary, in this section, we introduced an extensive examination of our proposed SA and AS cases of a kitten state, examined their phase space, and argued that such states offer sub-Planck structures that differ from a compass state in certain cases. Also it is suggested that the sensitivity in both SA and AS scenarios is significantly lower than the norm. In the next section, we present a full analysis of these results in relation to previous results, including brief physical explanations, implications, and beyond.
IV Discussion and remarks
@@
We now provide an overview of our findings by discussing their physical consequences and comparing them to previous research in the subject. This comprehensive discussion strives to integrate our findings into current understanding, providing a detailed picture of their effects and contributions to the field.
@@
The original compass state considered as a main example in our study [17], and our investigations and discussion are centered around this distinctive catlike state, which serves as the foundation for our exploration and examination. The compass state exhibits fascinating sub-Planck scale structures and enhanced sensitivity. However, these characteristics are lost when transitioning to a kitten state [34], which is essentially a smaller version of the cat state. This transition occurs when the macroscopic parameter is reduced, resulting in a shift from a cat state to a kitten state. Specifically, a bigger compass state exhibits pronounced sub-Planck features, indicating more refined quantum characteristics at scales lower than the Planck length; as discussed in §II.2 and then depicted in Fig. 1. Furthermore, the mean photon number in a bigger compass state is higher, indicating enhanced total photon content and intensity. This comparison demonstrates the significant differences in quantum behavior and measurement precision between the larger compass and kitten states; as observed in Fig. 3.
@@
Photon addition and subtraction operations on squeezed-vacuum states are extremely useful approaches for creating larger cat states [32, 33, 34, 35]. These methods have been demonstrated experimentally and provide an effective way to generate cat states of larger amplitude [34, 35]. In this study, we utilized the kitten version of the compass state and applied photon addition and subtraction operations with different order and magnitudes to construct our novel variants, as presented in Fig. 7. We then investigated the phase-space characteristics of these variants to gain insights into their quantum properties. The kitten version of the compass state, as shown in Fig. 1(a), which does not possess sub-Planckness, is now transformed into the states holding sub-Planck structures and demonstrating an enhanced sensitivity, implying the effectiveness of these multi-photon processes, as evident in Fig. 8 and Fig. 9. When the number of photons added increases, nonclassical features contained by these states are improved, but increasing photon subtraction destroys nonclassical structures in phase space. Note that this only occurs when the photon subtraction operation is combined with addition; otherwise, when photon subtraction is applied to a coherent state directly, the phase space remains unchanged [see Fig. 4]. Furthermore, photon addition raises the average photon count in the states, whereas photon subtraction maintains the photon statistics invariant.
This indicates that our states exhibit comparable phase-space attributes as the compass state when there are increased number of photon addition is performed. The photon subtraction plays an opposite role but we observed in our case their presence is also important. That is, in some cases the sub-Planck structures exhibited by our variants have isotropic form, as illustrated in Figs. 8(e) and 8(f). Furthermore, these variants have demonstrated a similar isotropic enhancement in sensitivity, as shown in Figs. 9(e) and 9(f). Except that with an appropriate photon subtractions the phase-space characteristics of our quantum states can also be controlled. This implies that our variants are an appropriate substitute to compass states, may be closely related to the current experiments, and can perform better under optimum conditions. Compasslike states built by superposition of compass states have also built isotropic forms of the sub-Planck structures [51], but our variants are obtained and followed a complete different technique, and presenting a completely different form in which photon addition and subtraction plays the main role in occurring of these intriguing characteristics. Also, note that the intricate superposition of higher number of quantum states can also be difficult to achieve in experiments. For example, as demonstrated in [14], coherent-state superpositions can be generated via third-order Kerr nonlinearity, which is relatively challenging to accomplish in experiments. Furthermore, catlike states are highly susceptible to loss, and because absorption cannot be ignored in currently available Kerr media, the capacity to extract coherent-state superposition before they fragile is severely limited [63], hence, comparatively, adding and subtracting photons from a quantum state is a better technique to construct such states [34, 35].
In conclusion, we presented an alternative versions of a compass state that exhibits sub-Planck structures and has an increased sensitivity than quasi-classical states (such as coherent states), making it suitable for quantum metrology applications. Our two variations achieved distinct properties, and in certain situations they perform better as compared to a compass state. Hence, our quantum states could serve as a reliable alternatives for these compass states. Further research is required to explore how to create these variants, which will require a novel and thorough investigation to develop new techniques for their generation. This endeavor will involve formulating innovative strategies and methodologies specifically designed to produce these advanced quantum states.
Acknowledgement

### v2

@@ IV Enhancement in sensitivity
Sensitivity and its relationship to the phase-space characteristics of a quantum state are thoroughly discussed in the §II.1. This concept is then applied to the compass state in the §II.2, where it is demonstrated that the presence of sub-Planck structures in those states significantly have enhanced their sensitivity to displacement and that optimizing controlling parameters may further enhance this sensitivity far better than the standard quantum limit. In §III.4, we thoroughly examined the phase space of the proposed SA and AS cases, confirming the presence of sub-Planck structures in their phase spaces. We now examine how these sub-Planck structures have impact on the sensitivity to phase-space displacement, which is analyzed by assessing the sensitivities using Eq. (3).
@@
$O_{\ket{\square_{i}}\bra{\square_{j}}}(\delta):=$ $G_{\alpha_{i},\alpha_{j}}\Lambda\sum^{r}_{n=0}(-1)^{n}\Gamma H_{r-n,q}\left[\text{i}\big{(}\alpha^{*}_{i}-\delta^{*}\big{)},\text{i}\alpha_{j}\right]$
×$\times H_{r-n,q}\left[-\text{i}\big{(}\alpha_{j}+\delta\big{)},-\text{i}\alpha^{*}_{i}\right]$ (35)
@@
$O_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\delta):=$ $(\alpha_{i}\alpha^{*}_{j})^{q}G_{\alpha_{i},\alpha_{j}}\Lambda\sum^{r}_{n=0}(-1)^{n}\Gamma(\alpha^{*}_{i}-\delta^{*})^{r-n}$
×$\times(\alpha_{j}+\delta)^{r-n}.$ (38)
@@
The corresponding sensitivities $S_{\ket{\square}}(\delta)$ and $S_{\ket{\blacksquare}}(\delta)$ with $\delta:=\nicefrac{{(\delta x+\text{i}\delta p)}}{{\sqrt{2}}}$ are shown in Fig. 11.
@@
Let us now examine the SA cases presented in the Figs. 11(a)-(c) belong to the SA situation, where it is evident that the overlap is zero for values $|\delta|<1$ (less than a coherent state) along arbitrary directions in phase space. This indicates that the sensitivity to displacement in this scenario surpasses the standard limit. The enhancement in sensitivity becomes more pronounced as the parameter $r$, which represents the number of added photons to the kitten state, increases. As shown in Figs. 11(a) and 11(b), this effect becomes clearly noticeable with a higher number of added photons. Specifically, the central structure is significantly reduced when the number of added photons increases from $r=12$ to $r=24$ with the number of subtracted photons, $q=12$, remains constant in this situation. Furthermore, the impact of increasing the number of eliminated photons from $q=12$ to $q=20$ is illustrated in the Figs. 11(b) and 11(c), where an enlargement in the central structure occurs, indicating that a higher value $|\delta|$ relative to the earlier case represented in Fig. 11(b) is needed to make the overlap zero. This suggests that the sensitivity to displacement in this circumstance decreases with an increment in $q$; hence, contrary to the parameter $r$, increasing the number of subtracted photons reduces sensitivity in the SA scenario.
@@
In the AS scenario, as depicted in Figs. 11(d)-(f), comparable to the SA example, an improved sensitivity to phase space displacements $\delta$ is observed. The values $|\delta|<1$ can render the $S_{\ket{\blacksquare}}(\delta)$ zero as observable by the central structure. Similar to the SA example, this amplification becomes more noticeable as $r$ increases, as illustrated in the instances depicted in Figs. 11(d) and 11(e), where an increment in the $r$ as shown in Fig. 11(e) has reduced the central structure corresponding to the overlap function, meaning that now a smaller value of $|\delta|$ is required to make the overlap zero compared to the case shown in Fig. 11(d). Furthermore, we observe that in Fig. 11(e) and Fig. 11(f), the photon subtraction $q$ also does not affect the sensitivity enhancement as depreciated; in Fig. 11(f), raising $q$ results in the same overlap as its preceding instance. However, the photon subtraction $q$ may have an effect on the isotropic nature of the sensitivity, as optimal parameter selections may produce isotropic regions, such as the circle-type regions centered at the origin observed in Figs. 11(g) to 11(i) show that in the AS scenario, sensitivity is consistently increased in all directions, making these instances better compared to their counter parts of SA and compass states.
V Outlook
@@
We now provide a brief discussion on our results with their summarized physical consequences. This comprehensive discussion strives to integrate our findings into current understanding, providing a detailed picture of their effects and contributions to the field.
@@
A four-headed kitten state is considered a main example in our study, which represents a smaller version of a compass state [13], and our investigations and discussion are centered around this state, which serves as the foundation for our exploration and examination. The compass state exhibits fascinating sub-Planck scale structures and enhanced sensitivity. However, these characteristics are lost when transitioning to a kitten state [52], which is essentially a smaller version of the cat state, as the one we presented in Fig. 1(a). This transition occurs when the macroscopic parameter is reduced, resulting in a transformation from a cat state to a kitten state. Specifically, a bigger compass state exhibits pronounced sub-Planck features, indicating more finer quantum characteristics at scales lower than the Planck length, as discussed in §II.2 and then depicted in Fig. 1. Note that the mean photon number in bigger catlike states is higher [24], indicating enhanced total photon content and intensity. This comparison demonstrates the significant differences in quantum behavior and measurement precision between the larger compass and kitten states, as observed in Fig. 3.
@@
Photon addition and subtraction operations on squeezed-vacuum states are extremely useful approaches for creating larger cat states [50, 51, 52, 53]. These methods have also been demonstrated experimentally and provide an effective way to generate cat states of larger amplitude [52, 53]. In this study, we utilized the kitten version of the compass state and applied photon addition and subtraction operations with different order and magnitudes to construct our novel variants, as presented in Figs. 8, 9 and 10. We then investigated the phase-space characteristics of these variants to gain insights into their quantum properties. The kitten version of the compass state, as shown in Fig. 1(a), which does not possess sub-Planckness, is now transformed into the states holding sub-Planck structures and demonstrating an enhanced sensitivity, implying the effectiveness of these multi-photon processes, as evident in Fig. 11.
When the number of photons added increases, nonclassical features contained by these states are improved, but increasing photon subtraction destroys nonclassical sub-Planck structures in phase space. Note that this only occurs for our SA case; otherwise, when photon subtraction is applied to a state directly (AS cases), the phase space nearly remains unchanged [see Fig. 5], and the size of the sub-Planck structures of the AS cases stays constant over the variation of the photon subtraction operations. Furthermore, photon addition raises the average photon count in the states, whereas photon subtraction maintains the photon statistics invariant. This demonstrates that the sequence of photon addition and subtraction operations in our case has a direct effect on the amount of quanta present in the states, which may be useful in maintaining the photon number in our states rather efficiently.
Our quantum states exhibit comparable phase-space attributes as the compass state when there is a large number of the photon addition of the photons is applied to the four-headed kitten state, and the role of the photon subtraction in our instances is also interesting as in some cases it slightly enhances the isotropy of these features. Distinct phase-space characteristics between our SA and AS examples are highlighted, namely that AS cases have smaller sub-Planck structures and so achieve higher sensitivity when compared to their SA counterparts. Our investigation implies that variants of the four-headed kitten state we provided are an appropriate substitute for compass states and may perform better compared to the compass state under optimum conditions.
VI Conclusion
We introduced alternative versions of the compass state, which are obtained by adding (and subtracting) an immense number of photons to a four-headed kitten state, with the option that the order in which these photons are applied to the state also changes and the noncommutativity of the bosonic operators results in two different quantum states with distinct phase-space features. Our outcomes revealed that the multiphoton operations we performed on the four-headed kitten state are quite effective and that these operations have transformed this multicomponent kitten state to other forms of states, which are richer in their nonclassical phase-space features and also exhibit sub-Planck structures. This investigation is relevant to previous experimental studies [52, 53], which demonstrated that photon operations are an effective means of enhancing the phase-space characteristics of quantum states, thereby supporting our findings.
The presence of crucial sub-Planck structures in the present investigation is influenced by the number of photons added or subtracted. Adding photons helps preserve these structures in both scenarios we explored, while subtracting photons typically disrupts them when photon addition is followed by photon subtraction. However, photon subtraction alone has no effect on the sub-Planck structures if applied before photon addition, and in this case (subtraction then addition), specific choices of photon subtraction operations may lead to isotropic versions of sub-Planck structures. These results directly apply to the sensitivity of these states as well. The sensitivity to displacement of the quantum states we proposed exceeds the standard limits, and this enhancement is also controlled by multiphoton operations. Specifically, increasing photon addition enhances sensitivity, while photon subtraction reduces it, in the sequence where photon addition is followed by subtraction. However, in the converse case of photon operations, the sensitivity to displacement remains unchanged over the variation of the photon subtraction, although, interestingly, for this case, improvement in the isotropy of sensitivity is observed for certain photon subtraction choices. The induction of sub-Planck structures in our indicated situations connects them to recent techniques for the development of nonclassical traits in quantum states [50, 51, 52, 53]. Perhaps these techniques may also be applied in the development of the quantum states we provided. Future research may inquire how to create our proposed quantum states, which will require a novel and thorough investigation to develop new techniques for their generation. This endeavor may involve formulating innovative strategies and methodologies specifically designed to produce these advanced quantum states.
Acknowledgment